# AMPLIFY:Attention-based Mixup for Performance Improvement and Label Smoothing in Transformer

### v2

@@ 5.3. Visualization of Experimental Results
In the semantic relationship diagram of the word "it" shown in Figure 4, because the sentence initially defines "film", expresses dissatisfaction with a certain type of movie, and finally expresses a positive attitude towards "film", it can be inferred that the "it" appearing in the sentence refers to "film" based on the consistency of semantic logic and the way the sentence expresses emotions. However, common pre-trained text models such as BERT split the keyword "cartoonish" into two lower-granularity tokens "cartoon" and "#ish" in order to solve the out-of-vocabulary problem, which obviously led EmbedMix to not consider that "cartoonish" is a complete adjective used to describe a certain movie genre. On the other hand, AMPLIFY not only focuses on "film", but also extends attention to the grammar dependencies between subjects and predicates such as "#ish", and "provides".
For the sake of fairness, both Mixup methods used in this section to compare respectively copy the outputs of the hidden layers at the same position, and perform Mixup after shuffling. Other settings are consistent with those in the 4.3 section. In addition, all attention visualization patterns in this section are from (Vig, 2019) and (Wang et al., 2021).
5.4. Further Studies on Hidden Mixup
@@
Table 4. The effects of applying the AMPLIFY operation to hidden layers of different depths in the model on seven benchmark datasets. The Input layer, Middle layer, and Last layer correspond to the MHA layers in the 4th, 8th, and 12th blocks, respectively. The experimental settings are the same as in section 4.3, and the values in the table are the average accuracy (%) and corresponding variance after running three times with three different random seeds.
@@
Research (Clark et al., 2019) and (Vig, 2019) found that the lower blocks (BERT layers 1-4) of BERT-base-uncased mainly learn vocabulary, syntax, and semantic information in the text sequence, the middle blocks (BERT layers 5-8) focus more on syntax information, and the last few blocks (BERT layers 9-12) focus on abstract semantic information and context-related information. Therefore, using Mixup in the lower layers can enhance the model’s robustness and generalization ability to word-level features. In the lower layers, BERT learns basic language features and contextual relationships between vocabulary, and using Mixup can enhance BERT’s perception of local information in the input text. Using Mixup in the middle layers can enhance the model’s robustness and generalization ability to sentence-level features. In the middle layers, BERT learns long dependency relationships between higher-level syntactic structures and morphemes, and using Mixup can enhance BERT’s ability to resist noise and out-of-domain features in the sequence. In theory, using Mixup in the high layers of the model should have the best effect, because in the high layers, BERT learns global feature information of the text, using Mixup can enhance BERT’s ability to recognize noisy and out-of-domain samples in the dataset, and thus improve its performance on downstream tasks. In addition, these global feature information is to some extent universal for different downstream tasks, so using Mixup in the high layers can effectively improve the model’s generalization ability and reduce the risk of overfitting. In summary, this also provides us with a better reason to perform Mixup on all layers.
6. Conclusion and Future Work
@@
The benchmark datasets and models utilized in this study were previously introduced in Section 4.1. These datasets are openly accessible and can be retrieved through the links provided in that particular section.
Compliance with Ethical Standards
Conflict of Interest: The authors declare that they have no conflict of interest.
Ethical approval: This article does not contain any studies with human participants or animals performed by any of the authors.
Publication ethics: The study was conducted in accordance with the publication ethics guidelines of the journal in which it is being submitted.

### v3

@@ Results
In the semantic relationship diagram of the word ”it” shown in Figure 8, because the sentence initially defines ”film”, expresses dissatisfaction with a certain type of movie, and finally expresses a positive attitude towards ”film”, it can be inferred that the ”it” appearing in the sentence refers to ”film” based on the consistency of semantic logic and the way the sentence expresses emotions. However, common pre-trained text models such as BERT split the keyword ”cartoonish” into two lower-granularity tokens ”cartoon” and ”#ish” in order to solve the out-of-vocabulary problem, which obviously led EmbedMix to not consider that ”cartoonish” is a complete adjective used to describe a certain movie genre. On the other hand, AMPLIFY not only focuses on ”film”, but also extends attention to the grammar dependencies between subjects and predicates such as ”#ish”, and ”provides”.
For the sake of fairness, both Mixup methods used in this section to compare respectively copy the outputs of the hidden layers at the same position, and perform Mixup after shuffling. Other settings are consistent with those in the Experimental Settings section. In addition, all attention visualization patterns in this section are from Vig, (2019) and Wang et al., (2021).
Further Studies on Hidden Mixup
@@
Table 6: The effects of applying the AMPLIFY operation to hidden layers of different depths in the model on seven benchmark datasets. The Input layer, Middle layer, and Last layer correspond to the MHA layers in the 4th, 8th, and 12th blocks, respectively. The experimental settings are the same as in section Experimental Settings, and the values in the table are the average accuracy (%) and corresponding variance after running three times with three different random seeds.
@@
Research Clark et al., (2019) and Vig, (2019) found that the lower blocks (BERT layers 1-4) of BERT-base-uncased mainly learn vocabulary, syntax, and semantic information in the text sequence, the middle blocks (BERT layers 5-8) focus more on syntax information, and the last few blocks (BERT layers 9-12) focus on abstract semantic information and context-related information. Therefore, using Mixup in the lower layers can enhance the model’s robustness and generalization ability to word-level features. In the lower layers, BERT learns basic language features and contextual relationships between vocabulary, and using Mixup can enhance BERT’s perception of local information in the input text. Using Mixup in the middle layers can enhance the model’s robustness and generalization ability to sentence-level features. In the middle layers, BERT learns long dependency relationships between higher-level syntactic structures and morphemes, and using Mixup can enhance BERT’s ability to resist noise and out-of-domain features in the sequence. In theory, using Mixup in the high layers of the model should have the best effect, because in the high layers, BERT learns global feature information of the text, using Mixup can enhance BERT’s ability to recognize noisy and out-of-domain samples in the dataset, and thus improve its performance on downstream tasks. In addition, these global feature information is to some extent universal for different downstream tasks, so using Mixup in the high layers can effectively improve the model’s generalization ability and reduce the risk of overfitting. In summary, this also provides us with a better reason to perform Mixup on all layers.
AMPLIFY is effectively a process of sample simulation, as the Attention matrix reflects the distribution of the association degree between feature elements. Therefore, blending the Attention output is akin to sampling from this distribution, creating a new hypothetical distribution. This assumed distribution concurrently contains the feature distribution of the original sample and introduces a certain level of randomness. The model, learning this fictitious distribution with random disturbance, essentially accepts the smoothed labels in the process, which ultimately facilitates soft label learning. Under soft labels, the model would lower its dependency on singular strong features and boost the learning of global features. Through repetitive mixups, this process effectively applies a degree of Laplace smoothing to the model, enhancing its robustness. In a statistical sense, Laplace smoothing can minimize the variables in the program, alleviating the impact of the assumed distribution bias on the results. Hence, from the perspective of Bayesian inference, the AMPLIFY algorithm, via Attention mixup, implements a soft-label Bayesian model to bolster its generalization capabilities.
Conclusion and Future Work
@@
Competing Interests
The authors declare there are no competing interests.
Author Contributions
  • Leixin Yang Conceptualization and Methodology, Software, Writing - Original Draft, Writing - Review & Editing.
  • Yu Xiang Conceptualization, Writing - Review & Editing.
@@
These benchmark datasets are openly accessible and can be retrieved through the links provided in that particular section.